_ 
For the equation x - Öx = 0, perform the following:
Solve for all values of x that satisfies the equation.


Answer by AnlytcPhil(1806)

                      _ 
For the equation x - Öx = 0, perform the following:
Solve for all values of x that satisfies the equation.
           _
      x - Öx = 0

Isolate the radical term
                _
           x = Öx

Square both sides
                 _
        (x)² = (Öx)²

          x² = x

Get 0 on right by subtracting x from both sides

      x² - x = 0

    x(x - 1) = 0

Use zero-factor principle

x = 0 is already solved

x - 1 = 0 gives x = 1

You should always check the solutions in equations containing
square roots because sometimes the solutions you get do not
check, but both these check.

So solutions are 0 and 1

Edwin
AnlytcPhil@aol.com